OpenStudy (anonymous):

-0.15x+0.1y+0.2z=0 0.1x-0.3y+0.2z=0 0.05x+0.2y-0.4z=0 x+y+z=1 Solve x, y , z

OpenStudy (anonymous):

If the decimal numbers are changed to fractions, then:\[\left\{x=\frac{16}{33},y=\frac{10}{33},z=\frac{7}{33}\right\}=\{0.485,0.303,0.212\} \]

OpenStudy (anonymous):

how did u solve it? can you telll me?

OpenStudy (anonymous):

I used Mathematica 8.04. Refer to the attachment.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

For your information:

OpenStudy (anonymous):

Thank you for the medal.